mgardner:

Solve the two-step equation. -9x + 0.4 = 4 Which operation must be performed to move all the constants to the right side of the equation? Then, which operation must be performed to isolate the variable? The solution to the equation is x = _____.

jhonyy9:

how you eliminate on the left side the +0,4 ? what you need adding to 0,4 to get zero ?

mgardner:

wait, what?

jhonyy9:

add to both sides -0,4 what will get so ?

jhonyy9:

or how you like subtract from both sides 0,4

mgardner:

you want me to add 0.4 to -9x and 4?

jhonyy9:

to add minus 0,4 to the both sides

mgardner:

oh ok

mgardner:

-9x + -0.4 = -9.4x 4 + -0.4 = 3.6

supie:

\(-9x + 0.4 = 4 \) What I would do is change the plus `(+)` in `+0.4` to a minus so it's `-0.4` so now that it's a minus subtract from both sides which should give you \(−9x=3.6\) because \(−9x+0.4−0.4=4−0.4\) \(4-0.4=3.6\) Therefore, you have \(−9x=3.6\) Then subtract `-9` from both sides \(\LARGE\frac{-9x}{-9}=\frac{3.6}{-9}\) \(\cancel{\frac{-9x}{-9}}=\frac{3.6}{-9}\) (-9x/-9 cancels) So all you would have left is \(\frac{3.6}{-9}\) once that is simplified then you should have your answer \(\frac{3.6}{-9}=?\) x=? If that makes any sense 👀

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(-9x + 0.4 = 4 \) What I would do is change the plus `(+)` in `+0.4` to a minus so it's `-0.4` so now that it's a minus subtract from both sides which should give you \(−9x=3.6\) because \(−9x+0.4−0.4=4−0.4\) \(4-0.4=3.6\) Therefore, you have \(−9x=3.6\) Then subtract `-9` from both sides \(\LARGE\frac{-9x}{-9}=\frac{3.6}{-9}\) \(\cancel{\frac{-9x}{-9}}=\frac{3.6}{-9}\) (-9x/-9 cancels) So all you would have left is \(\frac{3.6}{-9}\) once that is simplified then you should have your answer \(\frac{3.6}{-9}=?\) x=? If that makes any sense 👀 \(\color{#0cbb34}{\text{End of Quote}}\) good job @supie like a math teacher

supie:

Thanx

mgardner:

so x would = 3.6/-9 because if -9x/-9 cancels then all we would have is x. i hope you get what i mean.

mgardner:

wait i think i got it wrong hold up.

supie:

You have to find what x equals-

supie:

Simplify \(\frac{3.6}{-9}\) to get what x equals.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Simplify \(\frac{3.6}{-9}\) to get what x equals. \(\color{#0cbb34}{\text{End of Quote}}\) Or \(3.6 \div -9\)

mgardner:

3.6/-9 = -0.4 right?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mgardner 3.6/-9 = -0.4 right? \(\color{#0cbb34}{\text{End of Quote}}\) Yes

mgardner:

so x = -0.4!

supie:

Yes, correct!

mgardner:

ok, thanks a lot!

supie:

np